There are two factories in a small town. Both of them emit carbon dioxide into the air. Factory 1 currently emits 120 tons per month, whereas factory 2 currently emits 160 tons per month. The technology of each factory is different, so their costs of reducing emissions are different as well. The tables below show the costs of reducing emissions in increments of 20 tons per month for each factory:

Factory 1
Total cost of reducing emissions by 20 tons/month $50
Total cost of reducing emissions by 40 tons/month $150
Total cost of reducing emissions by 60 tons/month $270
Total cost of reducing emissions by 80 tons/month $410
Total cost of reducing emissions by 100 tons/month $570
Factory 2
Total cost of reducing emissions by 20 tons/month $20
Total cost of reducing emissions by 40 tons/month $60
Total cost of reducing emissions by 60 tons/month $110
Total cost of reducing emissions by 80 tons/month $200
Total cost of reducing emissions by 100 tons/month $300



The existing technology does not allow for reductions in emissions beyond 100 tons/month. That is, the most each factory could reduce its emissions by is 100 tons/month.

Suppose the government in this town would like to cut monthly emissions to half of the current level. To do that, the government has decided to impose a tax for every 20 tons of pollution per month emitted by a factory. To achieve its desired goal (but not exceed the goal), the tax would have to be set between $____________ and $____________ for every 20 tons/month. (The first number should be the lower end of the tax, and the second number should be the higher end of the tax.)
Answer #1

Current Emissions

Factory 1 = 120 tons

Factory 2 = 160 tons

Total Emissions = 120 + 160 = 280 tons

Now Government want to reduce emission to 50%

i.e = 50% of 280 tons = 140 tons

Now For both,

Tons reduced per month Factory 1 Cost Marginal Cost factory 1 Factory 2 Cost Marginal Cost factory 1
20 tons 50 50 20 20
40 tons 150 100 60 40
60 tons 270 120 110 50
80 tons 410 140 200 90
100 tons 570 160 300 100

So MArginal Cost of 100 will reduce it by 140 tons - 40 from Factory1 and 100 from factory 2

i.e Tax of 100 pe 20 tons can reduce it by 140 tons

so minimum is 100

If we go beyind 120 than tons reduec will be 160 as FActory 1 will reduce by 60 tons

So Range = 100 - 120 dollars per 20 tons

100 and 120 excluding